OpenStudy (anonymous):

Graph the function by starting with the graph of the basic function and then using the techniques of shifting, compressing, stretching, and /or reflecting. Graph one transformation in each coordinate system. Label each graph. f(x)=-2(x+1)^2+4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I would use something like desmos https://www.desmos.com/calculator to graph

jimthompson5910 (jim_thompson5910):

type into the first line y = x^2 this is the parent function

OpenStudy (anonymous):

Okay!

jimthompson5910 (jim_thompson5910):

then you'll shift the graph 1 unit to the left to get y = (x+1)^2 after that, you will multiply the (x+1)^2 term with -2 to get y = -2(x+1)^2. This will reflect the graph over the x axis and vertically stretch the graph by a factor of 2

OpenStudy (anonymous):

I have that already on one coordinate plane.

OpenStudy (anonymous):

Would shifting be in another coordinate plane?

jimthompson5910 (jim_thompson5910):

finally, you will add 4 to get y = -2(x+1)^2+4 this shifts the graph 4 units up

jimthompson5910 (jim_thompson5910):

each time you do a transformation, you will have it on its own xy axis

OpenStudy (anonymous):

I do not understand how do I show this on the (x,y)

jimthompson5910 (jim_thompson5910):

this is what it looks like if you did all equations I mentioned above on the same xy axis

jimthompson5910 (jim_thompson5910):

imagine each of those parabolas gets its own xy coordinate system

OpenStudy (anonymous):

I understand but my teacher wants us to show the change meaning when we plug in -1, 0 and 1

OpenStudy (anonymous):

I did get what you got, on each coordinate plane but I have to be able to show the process.

jimthompson5910 (jim_thompson5910):

so he wants you to plug those x values into each equation?

OpenStudy (anonymous):

basically

jimthompson5910 (jim_thompson5910):

you can create a table using desmos click on the equation you want to work with then click the gear icon I suggest duplicating the graph, so do so, but it's not mandatory then turn that duplicate graph into a table

jimthompson5910 (jim_thompson5910):

you can add or subtract x values from the table or change the x values that are already there the table will update itself with each change or addition in x

OpenStudy (anonymous):

Thank you! Now i got it! Are you able to help me with three more questions?

jimthompson5910 (jim_thompson5910):

it depends how long they are, but we'll see

OpenStudy (anonymous):

Okay, Suppose the point (2,4) is on the graph of y=f(x). Find a point on the graph of the given function . The reflection of the graph of y=f(x) across the y-axis

jimthompson5910 (jim_thompson5910):

Rule: reflecting any point (x,y) over the y axis will result in (-x,y)

OpenStudy (anonymous):

What does the question ask?

jimthompson5910 (jim_thompson5910):

it just wants you to reflect (2,4) over the y axis

OpenStudy (anonymous):

How would I do this?

OpenStudy (anonymous):

By graphing?

jimthompson5910 (jim_thompson5910):

by using that rule I posted

OpenStudy (anonymous):

So instead of 2,4 it would be -2,4

jimthompson5910 (jim_thompson5910):

(-2,4) don't forget the parenthesis

OpenStudy (anonymous):

Yes, and Would I graph this?

jimthompson5910 (jim_thompson5910):

you can if you want they just want you to find the point

OpenStudy (anonymous):

Okay , next question Solve the problem, If f(x)= x-3A/3x+5 and f(3)=6. what is the value of A?

OpenStudy (anonymous):

are you able to help me?